OpenStudy (anonymous):

tell me the correct method of finding the period of discrete time signal... the formula is N/r=2pi/w0; now having some confusions related to this

OpenStudy (gorv):

like f=sin2t w=2 period = 2pi/w0

OpenStudy (gorv):

\[N=\frac{ 2\pi }{ w_0 }\]

OpenStudy (gorv):

but period could be 2N 3N 4N 5N etc so we multiply by a number

OpenStudy (gorv):

called r

OpenStudy (gorv):

r=1,2,3,4,5,6,

OpenStudy (gorv):

\[N=r*\frac{ 2\pi }{ w_0 }\]

OpenStudy (gorv):

r=1

OpenStudy (gorv):

at r=1 N=2pi/w0

OpenStudy (gorv):

generally we take r=1 for fundamental period

OpenStudy (anonymous):

ri8.. but if N/r = 14/3.. then??

OpenStudy (anonymous):

The method of finding the Period of Discrete Time Signal is : \[N = \frac{2 \pi}{\omega_0} \times r\]

OpenStudy (anonymous):

Firstly evaluate \(\frac{\ 2\pi}{\omega_0}\), if its value is rational, then choose minimum value of \(r\) such that \(N\) comes to be Integer.. For example: If \(\frac{2 \pi}{\omega_0} = \frac{14}{3}\), you know \(\frac{14}{3}\) is rational, and how can you make it an Integer?? By choosing \(r\) as \(3\), such that \(3\) and \(3\) cancels, giving out \(14\) as Sample Period..

OpenStudy (anonymous):

And suppose \(\frac{2 \pi}{\omega_0} = \frac{3 \pi}{4}\), then it contains \(\pi\), which is an irrational value, so throw away this function, this function cannot be Periodic for any value of \(r\) you will choose.. :) Never..!!

OpenStudy (anonymous):

thanks..